# When Attention Collapses: How Degenerate Layers in LLMs Enable Smaller, Stronger Models

### v3

@@ A.3 Attention Pattern Visualization
A Potent (non-lazy) layer of a pre-trained GPT2 xLarge 48 layer model.
@@
A lazy layer of a pre-trained GPT2 xLarge 48 layer model.
Figure 13: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 xLarge model with 48 layers. The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30.
Appendix B Baselines
@@
Baselines trained from scratch (rand init.) :
@@
Baselines trained with various model initialization and efficient training techniques.
Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various zeroshot model initialization and effcient training techniques such as stacking, hybrid stacking, and half-width initialization. We explain these baseline training recipes using GPT-2 large and its variants as an example and apply the same process for other models.
@@
Appendix C Supplementary Experiments
We provide additional training data and plots for our main results discussed in Section 4.1 as shown in Table 3 and Figure 14. In Figure 5 (also refer Table 3 we compare our GPT-2 variants with baseline models trained form scratch. In Figure 14 (also refer Table 1) we compare our GPT-2 variants with baseline models trained using baseline zero-shot model initialization (and also re-training) techniques.
In Figure 15, we present the training curves of models trained during ablation as discussed in Section 4.2.
Knowledge Distillation
Recall we have already discussed distillation as a baseline in Section 4.1 and associated Figure 7. We perform an additional experiment in the same setting i.e. knowledge distillation as a baseline. Here we trained GPT-2 medium variants with 12 layers (half the number of a vanilla GPT-2 medium). We trained three models. First we distilled a 24-layer GPT-2 medium (teacher) to a 12-layer GPT-2 medium variant (student) and this student is initialized with all the alternate layers of the teacher. This setting is exactly same as discussed in DistillBERT Sanh et al. (2019). Next we trained two GPT-2 medium variants one from scratch (vanilla training) and the other with $\mathsf{Inheritune}$ recipe. Model trained with our recipe beats model trained with distillation. We defer a through investigation of distillation compared to $\mathsf{Inheritune}$ to future work.
How Inheritune addresses Attention Degeneration?
Recall we have discussed attention degeneration in Section 2 and attention patterns are visualized in Figure 12. Following up on our previous discussions in Figure 17 and Figure 18 we demonstrate that models trained with $\mathsf{Inheritune}$ has lesser lazy layers compared to it’s larger counterpart trained form scratch. We performed rank analysis for Figure 17 utilizing vanilla 24-layer GPT-2 medium and our 16-layer GPT-2 variant trained using $\mathsf{Inheritune}$. Additionally, we performed rank analysis for Figure 18 with a vanilla 48-layer GPT2 xLarge and a 24-layer GPT2 xLarge variant trained using $\mathsf{Inheritune}$.
Recall we have previously discussed that attention degeneration is connected with vanishing gradients of keys and queries Noci et al. (2022). The vanishing gradients is caused when the norm of the gradients Bengio et al. (1994) are so small that it fails to generate meaningful back-propagation signal. Since we are training smaller models intuitively $\lVert W_{Q}\rVert$ and $\lVert W_{K}\rVert$ should be smaller compared to their larger counterparts and hence the norm of gradients in the case of smaller models derived using $\mathsf{Inheritune}$ is higher leading to better training.
@@ Appendix C Supplementary Experiments
Table 3: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 xLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts. All models are trained for 100K steps, and reported using validation loss on the OpenWebText validation set. Note: GPT-2 Large and xLarge use one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds.
(a) GPT2-Medium
(b) GPT2-Large
(c) GPT2-xLarge
Figure 14: Models derived using Inheritune outperform three zero-shot initialization and efficient training baselines in terms of final validation loss. Our models demonstrate better convergence and generalization compared to all baselines. We trained GPT-2 xlarge, GPT-2 large and GPT-2 medium variants on OpenWebText-9B for 100K steps using baseline model initialization and efficient training techniques and our $\mathsf{Inheritune}$ training recipe.

### v4

@@ B.2 Attention Pattern Visualization of Potent and Lazy Layers
A Potent layer of vanilla GPT-2 XLarge.
@@
A lazy layer of vanilla GPT-2 XLarge.
Figure 12: Visualization of attention patterns in lazy and non-lazy layers of a vanilla GPT-2 XLarge model with 48 layers. The top row displays attention patterns for various heads (H) in layer (L) 8, while the bottom row shows patterns for layer (L) 30. For visual clarity, we display the full attention maps; however, attention in GPT-2 models is inherently causal.
Figure 13: Rank collapse worsens for larger LLMs, Inheritune helps to mitigate rank collapse. The maximum (max) rank across all attention heads for each layer is plotted, following the methodology in Figure 1. We analyze a 48-layer GPT-2 XLarge model which reveals rank-1 attention matrices in later layers (those beyond the halfway point), indicating rank collapse. Specifically, 22 out of the last 24 later layers exhibit rank-1 attention matrices (mean rank across all the 100 runs). Next, Our 24-layer GPT-2 XLarge variant, trained with $\mathsf{Inheritune}$, demonstrates improved rank across many layers, highlighting the effectiveness of our approach. Notably, 2 out of 12 of the later layers in our 24-layer variant exhibit rank-1 attention matrices.
Appendix C Understanding Attention Degradation using Attention Mass Analysis
In this paper, we have analyzed the attention degradation phenomenon primarily using a single metric-rank of the attention matrices (see Section 2). In this section, we aim to explore another thematically related metric to further investigate the nature of attention degradation.
In Figure 14, we present the layer-wise analysis of the attention matrix mass concentration in GPT-2 models. For this analysis, (similar to the rank analysis), we computed $A(X)$ using $N=100$ sequences selected at random from the validation set of OpenWebText (4.4M tokens), each with a sequence length of $T=100$ tokens across all attention heads within each layer. We define the average minimal column count $m$ required to capture 90% of the attention matrix mass for each head and layer as: $m^{(h,l)}=\frac{1}{N}\sum_{n=1}^{N}m^{*}_{n,h,l}$.Subsequently, we aggregate this metric per layer by taking the average across all heads: $\text{AvgMass}^{(l)}=\frac{1}{H}\sum_{h^{\prime}=1}^{H}m^{(h^{\prime},l)}$. We observe that many of the rank-collapsed attention matrices in deeper layers exhibit single-column attention structures, as measured by the $\text{AvgMass}^{(l)}$ criterion. As shown in Figure 15, we performed a mass analysis on contemporary billion-parameter OpenLLaMA models (Geng and Liu, 2023) and observed a similar pattern of attention degradation in the deeper layers. This provides concrete evidence that the phenomenon persists across a broad range of architectures and also at the billion-parameter scale.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XLarge
(a) OpenLLaMA-3B
(b) OpenLLaMA-7B
(c) OpenLLaMA-13B
Figure 15: The overall mass of attention matrices in billion-scale LLMs, pre-trained on trillions of tokens, tends to concentrate in fewer columns. This phenomenon becomes increasingly pronounced as the model size grows. We computed attention matrices using 100 tokens from a random subset of RedPajama with 1B tokens. Next, we performed 100 runs and plotted the mean and standard deviation of the mass as a function of layers for our mass analysis, respectively. We followed the same procedure as discussed in Section 2. Pre-trained checkpoints of OpenLLaMA-3B, OpenLLaMA-7B, and OpenLLaMA-13B (Geng and Liu, 2023), trained on 1T tokens from the RedPajama dataset Computer, 2023, were utilized. Overall, we observed that 90 % of the total mass of the attention matrices resides in fewer columns, with many attention matrices in the OpenLLaMA-13B model being single-column. This observation aligns closely with our analysis in Figure 1.
(a) Rank analysis of GPT-2 Medium
(b) Rank analysis of GPT-2 Large
(c) Rank analysis of GPT-2 XLarge
(d) Mass analysis of GPT-2 Medium
(e) Mass analysis of GPT-2 Large
(f) Mass analysis of GPT-2 XLarge
Figure 16: In standard decoder-style LLMs, attention matrices in deeper layers often degenerate into single-column matrices, leading to layers with fully degenerated attention that fail to learn meaningful representations. All models were trained on the OpenWebText dataset, and both rank and mass analyses were conducted using the FineWeb validation set, following the same procedure described in Figure 1. This further demonstrates the robustness of our analysis, as we reach the same conclusion using different evaluation datasets.
C.1 Data Robustness of Attention Degeneration Analysis
In Section 2 (Figure 1), we performed a rank analysis on three pre-trained GPT-2 models—Medium, Large, and XLarge using the validation set of the OpenWebText dataset, whose training split was originally used for pre-training these models. Here, we evaluate the data robustness of our analysis by repeating the same procedure on a validation set from FineWeb, a newer and distinct dataset. Except for the dataset substitution, all experimental steps remain identical to those described earlier. The results in Figure 16 consistently show that attention tends to lose rank, particularly in deeper layers, often collapsing into near single-column structures across all models. These findings further reinforce the robustness and generality of our observations. Moreover, for the LLaMA-3 model (Figure 11) and the OpenLLaMA models (Figure 15), we used publicly available model weights and conducted our analyses on off-the-shelf datasets that were not part of the models’ original training corpora.
Appendix D Baselines
@@
Baselines trained from scratch (random initializations) :
@@
Baselines trained with warm started training methods.
Here we compare our model derived using $\mathsf{Inheritune}$, to similar sized models trained with various model initialization and effcient training techniques which requires model to be initialized with trained weights such as stacking, hybrid stacking, and half-width. We explain these baseline training recipes using GPT-2 Large and its variants as an example and apply the same process for other models.
@@
Layers Initialization AvgRank Pre-train Val Loss ($\downarrow$)
4 rand N/A 3.25
4 1-4 layers from vanilla GPT2 8.40 3.22
4 5-8 layers from vanilla GPT2 9.48 3.19
4 9-12 layers (lazy layers) from GPT2 1.22 3.23
Table 4: Impact of initialization strategies on GPT-2 small variants. We analyzed the rank characteristics of a vanilla GPT2-small model (125M, 12 layers) trained on OpenWebText for 100K steps. Four-layer GPT2-small variants were initialized using the first 4 layers [1–4], middle 4 layers [5–8], last 4 layers [9–12], or with random initialization, and then trained for 100K steps on OpenWebText. Models initialized with the last 4 layers performed similarly to random initialization, while those initialized with layers exhibiting higher average max ranks achieved the best validation loss, regardless of proximity to the embedding layer. The training plots and rank analysis are provided in Figure 2.
Appendix E Supplementary Experiments
We provide supplementary tables and plots corresponding to the results discussed in the main paper, along with additional experiments, in this section. The final validation losses shown in Figure 2 are presented in Table 4. Similarly, the final validation losses from the training curves in Figure 5 and Figure 6 are summarized in Table 5. We also include training curves for models trained using various warm-started baselines (i.e., models initialized with learned weights) compared against our method, these results correspond to Table 2 and are visualized in Figure 18. Finally, Figure 19 presents the training curves for models used in the ablation study discussed in Section 4.3.
Distillation vs. Inheritune.
We conducted an additional experiment to compare $\mathsf{Inheritune}$ with knowledge distillation as a baseline. Specifically, we trained GPT-2 Medium variants with 16 layers under three different settings. First, we performed logit-based distillation Hinton et al. (2015), transferring knowledge from a 24-layer vanilla GPT-2 Medium (teacher) to a 16-layer student model. Second, we applied a DistilBERT-style distillation Sanh et al. (2019), where the student is initialized with the teacher’s layers. Finally, we trained a 16-layer GPT-2 Medium model from scratch using vanilla training. Across all comparisons, the model trained with our $\mathsf{Inheritune}$ recipe outperformed both distilled variants, achieving faster convergence and substantially better generalization after 50K training steps. We defer a thorough investigation of the relationship between $\mathsf{Inheritune}$ and distillation-based approaches to future work. The training configurations are provided in Section F.
@@ Appendix E Supplementary Experiments
Table 5: Inheritune achieves superior performance with reduced model size. Comparison of $\mathsf{Inheritune}$-trained models (24-layer GPT-2 XLarge, 18-layer GPT-2 Large, and 16-layer GPT-2 Medium) against full-sized counterparts and extended training baselines. The training steps of two different baselines are reported in the table, we use validation loss on the OpenWebText validation set. Note: GPT-2 Large and XLarge uses one round of $\mathsf{Inheritune}$; GPT-2 Medium uses three rounds. The lowest and our corresponding validation losses (lower is better) are highlighted in bold.
(a) GPT-2 Medium
(b) GPT-2 Large
(c) GPT-2 XLarge
Figure 18: Models derived using Inheritune outperform three warm-started baselines (Baseline-II) in terms of final validation loss. Our models demonstrate better convergence and generalization compared to all baselines. All the models are trained with OpenWebText for 100K steps. The curves are smoothed for visual clarity.